OpenStudy (anonymous):

Use the laws of exponents to simplify. 3^(3/4)*3^(7/8)

OpenStudy (anonymous):

that is the same as 3^0.75 * 3^0.875

OpenStudy (radar):

\[3^{3/4}\times3^{7/8}=3^{6/8}\times3^{7/8}=3^{13/8}=3\times3^{5/8}\]